OpenStudy (anonymous):

The volume of a fixed amount of gas at a fixed temperature varies inversely with pressure. If a gas has volume 4 L when the pressure is 900 N/m2, find its volume at a pressure of 1,200 N/m2.

OpenStudy (anonymous):

@mayankdevnani

OpenStudy (anonymous):

@97salonisharma

OpenStudy (anonymous):

\[P1V1 = P2V2\] Use this relation

OpenStudy (anonymous):

ie \[4*900 = 1200*V2\] Find V2

OpenStudy (anonymous):

hope this helps

OpenStudy (anonymous):

so would I do 3600+1200=V2?

OpenStudy (anonymous):

\[3600 = 1200*v2\] Divide 1200 on both sides

OpenStudy (anonymous):

\[V2 = 3L\]

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

Welcome...